OpenStudy (anonymous):

"Paraffin, a wax used to make candles, has a molecular formula of C25H52 (s) + 38 O2 ---> 23 CO2 (g) + 26 H2O (l) How many liters of Co2 are produced when a 23.4g paraffin candle burns at STP?"

OpenStudy (abb0t):

Do you know what STP means? standard temp & pressure and at STP, one mole of gas occupies 22.4 L of volume (molar volume) = \(\sf 22~ \frac{mol}{L}\)

OpenStudy (abb0t):

Take the mass given, conver to moles, and convert those moles of paraffin to CO\(\sf _2\) keeping in mind that that it's a 1:23 ratio for paraffin: CO\(\sf _2\) and finally, use STP to eliminate moles and get Liters.

OpenStudy (abb0t):

I will start you off: 23.5 g paraffin \(\times\) \(\sf \frac{1~mol}{353.77~ g ~parrafin}\) = __________ mol of paraffin \(\sf \times \frac{23~mol~CO_2}{1~mol~parrafin}\) ...and you can finish it off from here.

OpenStudy (anonymous):

Thanks for the help, abb0t! I'll try to figure it out. Don't have access to a calculator right now so idk if I can but thanks!

OpenStudy (anonymous):

@abb0t Ok, I used my phone calculator & got 1.53, now what do I do?

OpenStudy (abb0t):

You have moles of CO\(_2\) and you know at STP you have 22.7 mol/L....

OpenStudy (abb0t):

What is the relationship here.

OpenStudy (abb0t):

Your answer should be in the 10\(\sf ^{-2}\) values.

OpenStudy (eric_d):

34.24?

OpenStudy (abb0t):

Wrong.

OpenStudy (anonymous):

I'm lost.

OpenStudy (anonymous):

Do I divide 1.53/22.7? I'm sooo confused :(

OpenStudy (abb0t):

Ok, so you have 1.53 moles of CO\(_2\), and at STP, you have 22.7 moles/L it is the same as you did before, you cancel the units like in algebra: if you have: \(\sf \color{red}{\frac{a}{b} \times \frac{d}{a}}= \frac{d}{b}\) because the two \(\sf \color{red}{a}\)'s cancel out. Do you follow? It is the same with units: \(\sf \color{red}{23.4~grams ~paraffin} \times \frac{1~mol}{\color{red}{353.77~g~paraffin}}\) = mol

OpenStudy (abb0t):

\(\sf \color{red}{\cancel{23.4~grams ~paraffin}} \times \frac{1~mol}{\color{red}{{{\cancel{353.77~g~paraffin}}}}}\) = mol

OpenStudy (anonymous):

Yeah... still isn't making sense to me. >.<

OpenStudy (abb0t):

You're cancelling out the units. Because if you have: \(\sf \frac{grams~of~paraffin}{grams~of~paraffin} = 1 \) you're treating the units as constants or variables. If you have \(\sf \frac{4}{4} = 1\), same with \(\sf \color{red}{\frac{B}{B}} = 1\) AND SO FOrth.

OpenStudy (anonymous):

That's not why I'm confused. It's just hard for me to understand when you don't use the actual values :x

OpenStudy (anonymous):

I think I'm gonna take a break and come back to this problem or something..

OpenStudy (abb0t):

\(\sf \color{red}{23.4~grams ~paraffin} \times \frac{1~mol~paraffin}{\color{red}{353.77~g~paraffin} }= \frac{\color{red}{23.4~g~paraffin}(1~mol~paraffin)}{\color{red}{353.77~g~paraffin}} = \frac{23.4~mol}{353.77 }\) = 0.0686 mol of paraffin

OpenStudy (anonymous):

@abb0t so is .068 the answer? Cause that's what I got earlier but I thought I was wrong cause you said 10^-2 or something like that.

OpenStudy (eric_d):

It is

OpenStudy (anonymous):

No wait, it can't be cause I need Liters for the answer.

OpenStudy (eric_d):

Then it should be 34.24

OpenStudy (eric_d):

rite

OpenStudy (anonymous):

He said no so idk I'm lost.

OpenStudy (eric_d):

To get volume of one mole of any gas at STP , v got to multiply the moles by 22.4

OpenStudy (anonymous):

ok I'll write that down then. thanks love :)